if a sample of gas at sea level were moved to an elevation at which the pressure changed to 912 mm Hg, what would be the new elevation of the sample in kilometers
1
Expert's answer
2020-04-06T12:01:45-0400

Solution:

"760 mmHg = 10^5 Pa"

"912 mm Hg = 1.2 \times 10^5 Pa"


"P = hdg"

"1.2\times10^5 = h \times 1.225 \times 9.8"

h = 10 km


Answer: 10 km